OpenStudy (anonymous):

What is the slope of the line that passes through (-2, 7) and (4, 9)?

OpenStudy (anonymous):

Use the slope formula \[\frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\]

OpenStudy (anonymous):

(y2 - y1)/(x2-x1)

OpenStudy (anonymous):

m=1/3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Which of the following represents x =1/3 y written in general form?

OpenStudy (anonymous):

can you fan and give me best respons

OpenStudy (anonymous):

need help again

OpenStudy (anonymous):

what with?